# Rethinking Irregular Time Series Forecasting: A Simple yet Effective Baseline

### v3

@@ 4.5 Scalability and Efficiency Analysis
To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (TimeCHEAT (Liu, Cao, and Chen 2025), CRU (Schirmer et al. 2022), T-PatchGNN (Zhang et al. 2024)) on the PhysioNet dataset, with the batch size uniformly set to 32—see Figure 4. The results demonstrate that APN exhibits significant advantages across all key efficiency metrics. This is attributed to its core TAPA module, which efficiently generates information-condensed and regularized initial representations. These representations support a highly streamlined information aggregation and prediction architecture, ultimately achieving a lightweight and highly efficient model design.
@@ 5 Conclusion
This paper addresses the challenges of IMTS forecasting by proposing a general and efficient framework, APN. The core of APN lies in its novel Time-Aware Patch Aggregation (TAPA) module. This module introduces an aggregation-based paradigm, distinct from methods that rely on fixed patching or interpolation. By learning dynamic boundaries and applying a time-aware soft-weighting strategy, TAPA directly aggregates information from all raw observations to generate high-quality, regularized patch representations. This design choice not only ensures full data coverage and robustly handles information density variations, but also enables a streamlined and efficient overall architecture. Leveraging these superior representations, APN employs a concise query module and a shallow MLP to make predictions. Extensive experiments on multiple public IMTS benchmark datasets demonstrate that APN significantly outperforms existing state-of-the-art methods in both prediction accuracy and computational efficiency.

### v4

@@ 4.5 Scalability and Efficiency Analysis
To comprehensively evaluate the potential of APN in practical deployment, we conduct a comparative analysis of computational efficiency and resource consumption between APN and three representative baseline models (GraFITi (GraFITi), CRU (CRU), tPatchGNN (tPatchGNN)) on the USHCN dataset, with the batch size uniformly set to 32—see Figure 4. The results demonstrate that APN exhibits significant advantages across all key efficiency metrics. This is attributed to its core TAPA module, which efficiently generates information-condensed and regularized initial representations. These representations support a highly streamlined information aggregation and prediction architecture, ultimately achieving a lightweight and highly efficient model design.
@@ 5 Conclusion
This paper addresses the challenges of IMTS forecasting by proposing a general and efficient framework, APN. The core of APN lies in its novel Time-Aware Patch Aggregation (TAPA) module. This module introduces an aggregation-based paradigm, distinct from methods that rely on fixed patching or interpolation. By learning dynamic boundaries and applying a time-aware soft-weighting strategy, TAPA directly aggregates information from all raw observations to generate high-quality, regularized patch representations. This design choice not only ensures full data coverage and robustly handles information density variations but also enables a streamlined and efficient overall architecture. Leveraging these superior representations, APN employs a concise query module and a shallow MLP to make predictions. Extensive experiments on multiple public IMTS benchmark datasets demonstrate that APN significantly outperforms existing state-of-the-art methods in both prediction accuracy and computational efficiency.
6 Acknowledgments
This work was partially supported by the National Natural Science Foundation of China (No.62472174) and the Fundamental Research Funds for the Central Universities.